How many kg of water must be added to 50 kg of 80% acid to obtain a 60% acid?
Answer by stanbon(75887)

How many kg of water must be added to 50 kg of 80% acid to obtain a 60% acid?
-----
Equation:
water + water = water
----
0.20*50 + 1*x = 0.40(50+x)
----
20*50 + 100x = 40*50 + 40x
-----
60x = 20*50
x = (1/3)50
---
x = 16 2/3 kg (amt. of water to add)
======================================
Cheers,
Stan H.
===============